A boat is moving through a river at 9 km/hr. The river has a current of 2 km/hr. Assuming the current is either opposed to or in the direction of the movement of the boat, what is the maximum velocity of the boat over the riverbed?, what is the minimum? Assuming the current velocity is opposed to the velocity of the boat, to what speed should the boat slow so that its velocity is zero with respect to the river bed?
Answer by richwmiller(17219)

max=9+2=11
min=9-2=7
slow to 2 kmh to appear zero